OpenStudy (anonymous):

Multi-Step equations Need help on the one's I got wrong. 5(x - 4) = 12 for this I got 42/5 but it was marked wrong 7(2x + 6) = 1 I got 41/14 for this aswell,but marked wrong aswell 4(3 - 2x) = 15 and I got 12/23 for this Were did I go wrong???

OpenStudy (dobby1):

ok I am going to break this down for you so you can learn what to do

OpenStudy (anonymous):

Im all ears

OpenStudy (dobby1):

for the first one 5(x-4)=12 the fisrt step is to multiply 5 into the parenthases what do you get

OpenStudy (anonymous):

5x-20=12

OpenStudy (anonymous):

then I add +20 and add 20 to 12

OpenStudy (anonymous):

and 20 and 20 cancel each other out

OpenStudy (dobby1):

yes good

OpenStudy (dobby1):

what is 12 + 20

OpenStudy (anonymous):

the answer is 32/5 then?

OpenStudy (dobby1):

yes

OpenStudy (anonymous):

Now that is weird

OpenStudy (dobby1):

why

OpenStudy (anonymous):

I had 32/5 before and it was marked partiallly right

OpenStudy (anonymous):

the first two are marked partially right

OpenStudy (dobby1):

hm is this like a workpad or something?

OpenStudy (anonymous):

its online classes

OpenStudy (dobby1):

ok is this an essay?

OpenStudy (anonymous):

the 2nd one should be 41/14 but its being marked wrong no its just an assignment

OpenStudy (dobby1):

oh ok

OpenStudy (anonymous):

I really dont understand this for the 2nd and 3rd I keep getting the same answers and 3rd one is actually marked wrong

OpenStudy (dobby1):

ok the second one is a little tricky so I am going to have to do this one with you as well

OpenStudy (anonymous):

14x+42=1

OpenStudy (dobby1):

this is your question the first step is to multiply 7 into the parenthases 7(2x + 6) = 1

OpenStudy (dobby1):

yes

OpenStudy (dobby1):

14x+42=1 is right

OpenStudy (anonymous):

14/40?

OpenStudy (anonymous):

or 41

OpenStudy (dobby1):

ok lets subtract 42 from both sides what is 1 - 42 hint: it is 42-1 with a nagative sign next to it

OpenStudy (anonymous):

-41

OpenStudy (dobby1):

yes

OpenStudy (dobby1):

so now we have 14x=-41

OpenStudy (dobby1):

when you divide what do you get

OpenStudy (anonymous):

-2.9

OpenStudy (dobby1):

yeah I think you are only getting partial credit because you are sapposed to turn the fraction into a decimal so that is what you are doing wrong except for 2 and 3

OpenStudy (anonymous):

Hold on I'll see you a picture of it

OpenStudy (anonymous):

show you*

OpenStudy (dobby1):

ok

OpenStudy (anonymous):

do you see it

OpenStudy (dobby1):

yep

OpenStudy (anonymous):

confusing x.x

OpenStudy (dobby1):

yeah it is hm i dont know what they want

OpenStudy (dobby1):

i would try putting the anser we just got -2.9

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

yea your right

OpenStudy (dobby1):

oh good ok do you want helo with the third one?